Questions 48-49: You are conducting a randomized block design, testing a new type of medicine that’s supposed to lower cholesterol, with the following partial summary:

Source 

   SS

DF

Medicine

  5.20

  4

Block

  ___

__

Error

  0.54

  6

Total

12.91

19



48. What was the total sample size?: *

(A) 4

(B) 6

(C) 9

(D) 20

49. Find the block sum of squares: *

(A) 1.17

(B) 1.73

(C) 7.17

(D) 7.71

50. Which of the following are assumptions for the ANOVA procedure?

(I) Independent subjects
(II) Normality
(III) Equal variances
(IV) Nonnegative Data: *

(A) I, II, and III only

(B) II and III only

(C) II, III, and IV only

(D) I, II, III, and IV

(E) I, III, and IV only
Answer #1

48)

We know that,

df(Total) = n-1

So, 19 = n-1

n=20 which is the total sample size.

Hence option D is the correct option.

49)

We know that,

SS(Block) = SS(Total)-SS(Medicine)-SS(Error)

= 12.91-5.20-0.54

= 7.17

Hence option C is the correct option.

50)

We know that Independece, normality and equal vairances are the assumptions of the ANOVA procedure whereas Non-negative data is not.

Hence option A is the correct option.